Allen Products​ LP, wants to do a scenario analysis for the coming year. The pessimistic prediction for sales is $ 901,000​; the most likely amount of sales is $ 1,126,000​; and the optimistic prediction is $ 1,288,000. ​Allen's income statement for the most recent year is shown here

Allen Products, Inc. Income Statement for the Year Ended December 31, 2019     

Sales revenue   $937,200

Less: cost of goods sold 389,875

Gross profits      $547,325

Less: operating expenses             251,170

Operating profits              $296,155

Less: interest expense   32,802

Net profit before taxes $263,353

Less: taxes (rate 30%)     79,006

Net profits after taxes    $184,347

a. Use the ​percent-of-sales method, the income statement for December​ 31,2019​, and the sales revenue estimates to develop ​pessimistic, most​ likely, and optimistic pro forma income statements for the coming year.

b. Explain how this method could result in overstatement of profits for the pessimistic case and understatement of profits for the most likely and optimistic cases.

c. Restate the pro forma income statements prepared in part a. to incorporate the following assumptions about the costs:

$208,681 of the cost of goods sold is​ fixed; the rest is variable. $198,070 of the operating expenses is​ fixed; the rest is variable. All the interest expense is fixed.​

d. Compare your findings in part c. to your findings in part a. Do your observations confirm your explanation in part b​?

Use the ​percent-of-sales method, the income statement for December​ 31, 2019, and the sales revenue estimates to develop​pessimistic, most​ likely, and optimistic pro forma income statements for the coming year.

Complete the pro forma income statement for the year ending December​ 31, 2020 that is shown below​ (pessimistic scenario): ​(Round the percentage of sales to one decimal place and the pro forma income statement accounts to the nearest​ dollar.)
Answer #1
Year Ended December 31, 2019
A B=A/937200
Actual Percent of Sales Pessimistic Most Likely Optimistic
C Sales Revenue $937,200 100% $901,000 $1,126,000 $1,288,000
D Less: Cost of goods sold $389,875 42% $374,816 (42%*901000) $468,416 (42%*1126000) $535,808 (42%*1288000)
E=C-D Gross Profit $547,325 $526,184 $657,584 $752,192
F Less: Operating expenses $251,170 27% $241,468 (27%*901000) $301,768 (27%*1126000) $345,185 (27%*1288000)
G=E-F Operating Profit $296,155 $284,716 $355,816 $407,008
H Less: Interest Expense $32,802 4% $31,535 (4%*901000) $39,410 (4%*1126000) $45,080 (4%*1288000)
I=G-H Net Profit before taxes $263,353 $253,181 $316,406 $361,928
J=I*30% Less: taxes (30%) $79,006 $75,954 $94,922 $108,578
K=I-J Net Profit after taxes $184,347 $177,227 $221,484 $253,349
b This method assumes that Cost of goods sold, Operating Expenses and Interest expenses are variable and proportional to sales
Actually, part of these expenses may be fixed which do not change with Sales volume
Hence , the expenses are understated in pessimistic case and overstated in optimistic scenario
Thus , Profits are overstated in pessimistic scenario and understated in optimistic scenario
c Pessimistic Scenario:
Cost of goods sold $382,876 208681+(389875-208681)*(901000/937200)
Operating Expenses $249,119 198070+(251170-198070)*(901000/937200)
Most Likely Scenario
Cost of goods sold $426,377 208681+(389875-208681)*(1126000/937200)
Operating Expenses $261,867 198070+(251170-198070)*(1126000/937200)
Optimistic Scenario
Cost of goods sold $457,697 208681+(389875-208681)*(1288000/937200)
Operating Expenses $271,046 198070+(251170-198070)*(1288000/937200)
Actual 2019 Pessimistic Most Likely Optimistic
C Sales Revenue $937,200 $901,000 $1,126,000 $1,288,000
D Less: Cost of goods sold $389,875 $382,876 $426,377 $457,697
E=C-D Gross Profit $547,325 $518,124 $699,623 $830,303
F Less: Operating expenses $251,170 $249,119 $261,867 $271,046
G=E-F Operating Profit $296,155 $269,005 $437,756 $559,257
H Less: Interest Expense $32,802 $32,802 $32,802 $32,802
I=G-H Net Profit before taxes $263,353 $236,203 $404,954 $526,455
J=I*30% Less: taxes (30%) $79,006 $70,861 $121,486 $157,937
K=I-J Net Profit after taxes $184,347 $165,342 $283,468 $368,519